OpenStudy (anonymous):

Find the exact value of the expression 4sin(pi/3)-3cot(pi/3)

OpenStudy (anonymous):

= \[\sqrt{3}\]

OpenStudy (anonymous):

correct?

OpenStudy (mertsj):

\[4(\frac{\sqrt{3}}{2})-3(\frac{\sqrt{3}}{3}) = 2\sqrt{3}-\sqrt{3}=\sqrt{3}\]

OpenStudy (mertsj):

good job!!

OpenStudy (anonymous):

tahnks

OpenStudy (mertsj):

yw